Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
amagazzu - 16:57 Wednesday 31 May 2023 (60384) Print this report
Operator Report - Morning shift

Upon reaching the site I found the ITF locked in CARM_NULL_1F since 5:56 UTC. It unlocked at 6:46 UTC due to the ongoing DET activity to check the OMC behavior, carried out by Gouaty and Hui.
The ISC Crew (Mantovani, Spinicelli), with the support of the DET experts, were able to reach LOW_NOISE_1 at 8:56 UTC and LOW_NOISE_2 at 10:02 UTC. The ITF unlocked at 10:41 UTC for no clear reason.
We attempted to lock again, but the crew noticed some issues with the WI suspension behavior.
Balancing of the Suspensions, carried out by Ruggi, in progress.

Other activies carried out during the shift:
PA actuator characterization tests (see report #60382).

Sub-system reports

At 8:43 UTC, under request of Mantovani, I enabled the Etalon NI and WI loops. We disabled them again at 8:47 UTC due to the high signal error, it will be re-engaged when the temperature will decrease and get close to their setpoints.

Images attached to this report
Comments to this report:
tournefi - 9:52 Friday 02 June 2023 (60418) Print this report

The unlock reported by the operator in this entry, which occured for "no clear reason" (on May 31st at 10:41) happened within a few minutes when Maria finished the work in the TCS room.

Looking at at TCS microphone the coincidence is quite clear. 

Images attached to this comment
Search Help